40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 3 22:37:44 2012, Jenne, Update, General, OpLev 90-day trend OpLevTrends_90days_Ending3May2012
    Reply  Fri May 4 17:44:42 2012, Jenne, Update, General, OpLev 90-day trend 
Message ID: 6601     Entry time: Thu May 3 22:37:44 2012     Reply to this: 6602
Author: Jenne 
Type: Update 
Category: General 
Subject: OpLev 90-day trend 

After fixing the PRM tonight, all of the oplevs look okay. 

.....except ITMX, which looks like it's power dropped significantly after the CDS upgrade.  To be investigated tomorrow.

Attachment 1: OpLevTrends_90days_Ending3May2012  24 kB
ELOG V3.1.3-